An athlete runs at a rate of 3 meters per second around a 400- meter track. Which equation best represents the distance covered if he has already run 2 laps?
A. D=400 + 3s
B. D=800 + 3s
C. D=800 + 6s
D. D=800 - 3s

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
Let s=number of seconds that he runs after he has already ran 2 laps
Total Distance covered=D=2*400+3*s or
D=800 + 3s

Hope this helps---ptaylor